OpenStudy (anonymous):

Trying to remember how to factor 1600X^2+880xy+121y^2

OpenStudy (anonymous):

(40x + 11y)^2

OpenStudy (anonymous):

But how does that work out to the 880xy?

OpenStudy (kinggeorge):

Well, you can see that 1600 = 40^2 and 121 = 11^2, so the first guess you should make is \[(40x +11y)(40x+11y)\] and see if it works. In this case, it does.

OpenStudy (anonymous):

OK, looks like I was having a major mental block.

OpenStudy (kinggeorge):

If you FOIL it, the 880 just pops right out.

OpenStudy (anonymous):

Right, I was thinking 40*11 was 44

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

Thanks